OpenStudy (anonymous):

PLEASE HELP I have tried to work it out but just cant... What is the 32nd term of the arithmetic sequence where a1 = 13 and a13 = −59? −185 −179 −173 −167

jimthompson5910 (jim_thompson5910):

The nth term in general is an = a1 + d(n-1) where d is the common difference

OpenStudy (anonymous):

i think 6 is the common difference...

jimthompson5910 (jim_thompson5910):

a1 = 13 so, an = a1 + d(n-1) an = 13 + d(n-1)

OpenStudy (anonymous):

-6 i mean

jimthompson5910 (jim_thompson5910):

then we also know that a13 = -59 (ie the 13th term is -59) an = 13 + d(n-1) a13 = 13 + d(13-1) ... replace n with 13 -59 = 13 + d(13 - 1) solving that for d gives you d = -6, correct

OpenStudy (anonymous):

and then that is where i got stuck at...

jimthompson5910 (jim_thompson5910):

so that means an = a1 + d(n-1) an = 13 + d(n-1) an = 13 + (-6)(n-1) an = 13 - 6n + 6 an = -6n + 19

jimthompson5910 (jim_thompson5910):

an = -6n + 19 is the nth term formula that allows you to find any term you want the 32nd term, so plug in n = 32 and evaluate

OpenStudy (anonymous):

so a32+-6(32)+19=... right?

jimthompson5910 (jim_thompson5910):

a32 = -6(32) + 19 I think you meant

OpenStudy (anonymous):

yeah, my bad, i was pushing the shift key....

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

now just evaluate -6(32) + 19

OpenStudy (anonymous):

-211 is what i got... i did something wrong.

jimthompson5910 (jim_thompson5910):

You should have gotten -173

jimthompson5910 (jim_thompson5910):

alternatively, you can plug n = 32 into an = 13 + (-6)(n-1) to get an = 13 + (-6)(n-1) a32 = 13 + (-6)(32-1) a32 = -173

jimthompson5910 (jim_thompson5910):

hmm you must have done 6*32 + 19 and made the result negative

OpenStudy (anonymous):

i got lost at the part where its a32=13+(-6)(32-1)... and yeah, i thought thats what i was supposed to do because () usually indicates *...

OpenStudy (anonymous):

right? so what was i actually supposed to do for this equation? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

either equation works since an = 13 + (-6)(n-1) is equivalent to an = -6n + 19

OpenStudy (anonymous):

ohh, okay that makes much more since now!!

OpenStudy (anonymous):

so instead of using () for multiplication, i use it for addition?

jimthompson5910 (jim_thompson5910):

no it means multiplication (-6)(n-1) means -6 times n-1

jimthompson5910 (jim_thompson5910):

you may need to use the symbol * for multiplication

OpenStudy (anonymous):

BEHHHH I HATE MATH~

jimthompson5910 (jim_thompson5910):

yeah it's picky sometimes

OpenStudy (anonymous):

yes! VERY! i think i may understand it...

jimthompson5910 (jim_thompson5910):

that's great, glad you do